# Improving Adversarial Robustness via Phase and Amplitude-aware Prompting

### v1

@@ Abstract
Deep neural networks are found to be vulnerable to adversarial noises. The prompt-based defense has been increasingly studied due to its high efficiency. However, existing prompt-based defenses mainly exploited mixed prompt patterns, where critical patterns closely related to object semantics lack sufficient focus. The phase and amplitude spectra have been proven to be highly related to specific semantic patterns and crucial for robustness. To this end, in this paper, we propose a Phase and Amplitude-aware Prompting (PAP) defense. Specifically, we construct phase-level and amplitude-level prompts for each class, and adjust weights for prompting according to the model’s robust performance under these prompts during training. During testing, we select prompts for each image using its predicted label to obtain the prompted image, which is inputted to the model to get the final prediction. Experimental results demonstrate the effectiveness of our method.
@@ 1 Introduction
Motivated by the above studies, we propose a Phase and Amplitude-aware Prompting (PAP) defense mechanism, which constructs phase and amplitude prompts to stabilize the model’s predictions during testing.
@@ 2.1 Adversarial Attacks
Adversarial attacks craft malicious noises to mislead target models. White-box attacks like Projected Gradient Descent (PGD) [7], AutoAttack (AA) [25], Carlini&Wagner (C&W) [26] and Decoupling Direction and Norm (DDN) [27] craft noises through accessing and utilizing models’ intrinsic information like structures and parameters. For black-box attacks like transfer-based attacks and query-based attacks [28], attackers have no access to the models’ internal information, and thus attack only by interacting with model’s inputs and outputs.
@@ 3.2 Motivation
DNNs can be easily fooled by adversarial noises. Defenses like adversarial training and denoising methods all improve robustness with a high computational cost, promoting the development of prompt-based defenses due to the efficiency. However, existing prompt-based defenses focus on mixed patterns like pixel or frequency information, which cannot capture specific patterns like structures and textures. Thus, we seek to further disentangle these patterns for robustness.
@@ 3.3.1 Prompt Construction and Training
where $\tilde{x}_{i}^{p^{\prime}}$ is the prompted adversarial example using prompts from $y^{\prime}_{i}$, which is a randomly selected label different from its ground-truth label $y_{i}$. $h_{\theta}^{y^{\prime}_{i}}(\cdot)$ and $h_{\theta}^{y_{i}}(\cdot)$ denote outputs on $y^{\prime}_{i}$ and $y_{i}$, and $\tau$ is a threshold set as 0.1 in this paper.
@@
Table 1: The impact of different spectra on robustness. Adv. All denotes normal noises. Nat. Pha./Amp. indicate we replace phase/amplitude spectra with corresponding natural spectra.
@@ 3.3.2 Weighting Method
where we use Equation 7 for training, and we adjust the weight every 5 epochs for the efficiency. The finally learned weight $w^{*}$ is utilized for testing, and is shown in Table 5. The weighting strategy is illustrated in Figure 3.
@@ 4.1 Experimental Settings
Defense Settings. We use prompt-based defenses C-AVP [12] and Freq [11] as baselines, where they are designed only for defending on naturally pre-trained models. We use natural training (NAT), AT [7], TRADES [29] and MART [30] to obtain pre-trained models. We use PGD with perturbation budget 8/255, perturb step 10 and step size 2/255 for training. We train prompts by SGD [38] for 100 epochs, where the initial learning rate is 0.1 and is divided by 10 at the 75-th epoch. The batch size is 512 for CIFAR-10, and 256 for Tiny-ImageNet. We set $\lambda_{1}\!=\!3$, $\lambda_{2}\!\!=\!\!400$, $\lambda_{3}\!=\!4$ for naturally pre-trained models, and $\lambda_{1}\!=\!1$, $\lambda_{2}\!=\!5000$, $\lambda_{3}\!=\!4$ for adversarially pre-trained models. We omit deviations in several tables due to their small values ($\leq\!\!0.60\%$). More details can be found in Appendix C.
@@ 4.5 Ablation Studies
Comparison with Universal Prompts. To verify the superiority of learning prompts for each class, we train a universal phase-level prompt and a universal amplitude-level prompt for comparisons. Table 10 shows performances under universal prompts are worse than those of ours, indicating that our method helps enhance the robustness.
Effectiveness When Blurring the Edges. Some attacks like DDN tend to disrupt edges of objectives. Therefore, it’s natural to question whether robustness gains from our PAP come from edge blurring. To this end, we apply Gaussian Blur on the tested image for evaluations. As shown in Table 11, when blurring edges, our method can still achieve superior defenses, indicating the effectiveness of PAP does not come from edge blurring.
@@ 5 Conclusion
In this paper, we focus on specific semantic patterns for improving prompt-based defenses. It has been proven that phase and amplitude spectra reflect structures and textures, and both of them need to be manipulated for robustness. Therefore, we construct prompts using these spectra, and propose a Phase and Amplitude-aware Prompting (PAP) defense, which learns a phase-level prompt and an amplitude-level prompt for each class. Considering different influences of phase-level and amplitude-level prompts for robustness, we design a weighting method for them according to the robustness under these prompts. To perform testing efficiently, we select prompts according to predicted labels, and design a data-prompt mismatching loss to mitigate the negative effects of mismatches between images and their selected prompts. Experimental results demonstrate our method helps defend against general attacks and adaptive attacks, achieving superior transferability. The limitation of our work is that we sacrifice some natural accuracy when prompting on naturally pre-trained models. We will address it in the future such as using Contrastive Learning [39]. Overall, our defense explores specific semantic patterns to improve performances of prompt-based defenses.
@@ Appendix B Prompt Selection Method for Testing from C-AVP
where $p$ is the selected prompt for the test image $x_{test}$, while $p_{i}$ is the prompt of class $i$ and $h_{\theta}^{i}$ is the output of class $i$, and $\mathcal{C}$ denotes the set of classes. Clearly, when the number of class becomes large, this strategy for testing can easily cause extremely high computational costs. Also, when testing using it, C-AVP indeed achieves some robustness, but sacrifices natural accuracy a lot, which can be seen in Table 4.
@@ Appendix G Visualizations of Prompted Images
We present additional visualized results of prompted images using our prompts, which are presented as follows. It can be seen that our method retains complete and natural semantic patterns after prompting.

### v2

@@ Abstract
Deep neural networks are found to be vulnerable to adversarial perturbations. The prompt-based defense has been increasingly studied due to its high efficiency. However, existing prompt-based defenses mainly exploited mixed prompt patterns, where critical patterns closely related to object semantics lack sufficient focus. The phase and amplitude spectra have been proven to be highly related to specific semantic patterns and crucial for robustness. To this end, in this paper, we propose a Phase and Amplitude-aware Prompting (PAP) defense. Specifically, we construct phase-level and amplitude-level prompts for each class, and adjust weights for prompting according to the model’s robust performance under these prompts during training. During testing, we select prompts for each image using its predicted label to obtain the prompted image, which is inputted to the model to get the final prediction. Experimental results demonstrate the effectiveness of our method.
@@ 1 Introduction
Motivated by the above studies, we propose a Phase and Amplitude-aware Prompting (PAP) defense mechanism, which constructs phase and amplitude-level prompts to stabilize the model’s predictions during testing.
@@ 2.1 Adversarial Attacks
Adversarial attacks craft malicious noises to mislead target models. White-box attacks like Projected Gradient Descent (PGD) attack [7], AutoAttack (AA) [25], Carlini&Wagner (C&W) [26] and Decoupling Direction and Norm (DDN) [27] craft noises through accessing and utilizing models’ intrinsic information like structures and parameters. For black-box attacks like transfer-based attacks and query-based attacks [28], attackers have no access to the models’ internal information, and thus attack only by interacting with model’s inputs and outputs.
@@ 3.2 Motivation
DNNs can be easily fooled by adversarial noises. Defenses like adversarial training and denoising methods all improve robustness with a high computational cost, promoting the development of prompt-based defenses due to the efficiency. However, existing prompt-based defenses focus on mixed patterns like pixel or frequency information, which cannot capture specific patterns like structures and textures [32, 33]. Thus, we seek to further disentangle these patterns for robustness.
@@ 3.3.1 Prompt Construction and Training
where $\tilde{x}_{i}^{p^{\prime}}$ is the prompted adversarial example using prompts from $y^{\prime}_{i}$, which is a randomly selected label different from its ground-truth label $y_{i}$. $h_{\theta}^{y^{\prime}_{i}}(\cdot)$ and $h_{\theta}^{y_{i}}(\cdot)$ denote outputs on $y^{\prime}_{i}$ and $y_{i}$, and $\tau$ is a threshold.
@@
Table 1: The impact of different spectra on robustness. Adv. All denotes normal noises. Nat. Pha./Amp. indicates we replace phase/amplitude spectra with corresponding natural spectra.
@@ 3.3.2 Weighting Method
where we use Equation 7 for training. The finally learned weight $w^{*}$ is utilized for testing, and is shown in Table 5. The weighting strategy is illustrated in Figure 3.
@@ 4.1 Experimental Settings
Defense Settings. We use prompt-based defenses C-AVP [12] and Freq [11] as baselines, where they are designed only for defending on naturally pre-trained models. We use natural training (NAT), AT [7], TRADES [29] and MART [30] to obtain pre-trained models. We use PGD with perturbation budget 8/255, perturb step 10 and step size 2/255 for training. We train prompts by SGD [40] for 100 epochs, where the initial learning rate is 0.1 and is divided by 10 at the 75-th epoch. The batch size is 512 for CIFAR-10, and 256 for Tiny-ImageNet. We set $\lambda_{1}\!=\!3$, $\lambda_{2}\!\!=\!\!400$, $\lambda_{3}\!=\!4$ for naturally pre-trained models, and $\lambda_{1}\!=\!1$, $\lambda_{2}\!=\!5000$, $\lambda_{3}\!=\!4$ for adversarially pre-trained models. The threshold $\tau$ is set as 0.1, and we adjust the weights of amplitude-level prompts every 5 epochs. We omit deviations in several tables due to their small values ($\leq\!\!0.60\%$). More details can be found in Appendix C.
@@ 4.5 Ablation Studies
Comparison with Universal Prompts. To verify the superiority of learning prompts for each class, we train a universal phase-level prompt and a universal amplitude-level prompt for comparisons. For fairness, the data-prompt mismatching loss is removed on universal prompts, and other settings for universal prompts are the same as those from our method. Table 10 shows performances under universal prompts are worse than those of ours, indicating that our method helps enhance the robustness.
Effectiveness When Blurring the Edges. Some attacks like DDN tend to disrupt edges of objectives. Therefore, it’s natural to question whether robustness gains from our PAP come from edge blurring. To this end, we apply Gaussian Blur on the test image for evaluations. As shown in Table 11, when blurring edges, our method can still achieve superior defenses, indicating the effectiveness of PAP does not come from edge blurring.
5 Limitation
Despite the advances in adversarial defenses, our method still has several limitations. First, our method sacrifices some natural accuracy when prompting on naturally pre-trained models. We will address it in the future such as using Contrastive Learning, since it is a useful method for mitigating the trade-off problem between natural and robust accuracies [41, 42, 43]. Second, we do not perform evaluations on ImageNet due to the limited computational resources. However, we conduct experiments on Tiny-ImageNet which has been widely used. Tiny-ImageNet-200 is larger and has a larger resolution than CIFAR-10, with more numbers of classes than those of ImageNet-100. Results show that our method achieves superior performances, leading us to believe that our method can also work well on ImageNet. We leave them to the future work.
@@ 6 Conclusion
In this paper, we focus on specific semantic patterns for improving prompt-based defenses. It has been proven that phase and amplitude spectra reflect structures and textures, and both of them need to be manipulated for robustness. Therefore, we construct prompts using these spectra, and propose a Phase and Amplitude-aware Prompting (PAP) defense, which learns a phase-level prompt and an amplitude-level prompt for each class. Considering different influences of phase-level and amplitude-level prompts for robustness, we design a weighting method for them according to the robustness under these prompts. To perform testing efficiently, we select prompts according to predicted labels, and design a data-prompt mismatching loss to mitigate the negative effects of mismatches between images and their selected prompts. Experimental results demonstrate our method helps defend against general attacks and adaptive attacks, achieving superior transferability. Overall, our defense explores specific semantic patterns to improve performances of prompt-based defenses.
Acknowledgements
This work was supported in part by the National Natural Science Foundation of China under Grants U22A2096, 62036007 and 62306227, in part by Scientific and Technological Innovation Teams in Shaanxi Province under grant 2025RS-CXTD-011, in part by the Shaanxi Province Core Technology Research and Development Project under grant 2024QY2-GJHX-11, in part by the Fundamental Research Funds for the Central Universities under Grant QTZX23042, in part by the Innovation Fund of Xidian University under Grant YJSJ25007.
@@ Appendix B Prompt Selection Method for Testing from C-AVP
where $p$ is the selected prompt for the test image $x_{test}$, while $p_{i}$ is the prompt of class $i$ and $h_{\theta}^{i}$ is the output of class $i$, and $\mathcal{C}$ denotes the set of classes. Clearly, when the number of classes becomes large, this strategy for testing can easily cause extremely high computational costs. The prompt selection strategy of C-AVP is inefficient on numerous classes, and results in Table 4 show baselines with this strategy lose natural accuracy a lot. In comparison, our prompt selection strategy is efficient on numerous classes, and our defense with this strategy achieves superior defenses with higher natural accuracy, verifying the superiority of our prompt selection strategy.
@@ Appendix F Hyper-parameter Studies
There exists a trade-off problem in our method. As shown in Figure 5, for naturally pre-trained models, the natural accuracy increases while the robust accuracy drops as $\lambda_{1}$ or $\lambda_{2}$ increases. As shown in Figure 6, for adversarially pre-trained models, when $\lambda_{2}$ varies from 0 to 5000, the trade-off problem exists explicitly. Overall, the hyper-parameters we set achieve superior performances in both natural and robust accuracies.
@@ Appendix G Visualizations of Prompted Images
We present additional visualized results of prompted images using our prompts, which are presented as follows. Here, following previous works [45, 46, 47], C-AVP performs prompting in the pixel space by adding random noises to the surrounding area inside the image, only keeping the square area in the center unchanged. Therefore, C-AVP is only a frame. It can be seen that our method retains complete and natural semantic patterns after prompting.
C-AVP performs prompting by adding noises around the image in the pixel domain, while Freq performs prompting on the high-frequency domain. They both train their prompts without considering their disruptions on the natural semantic patterns. In comparison, our method construct prompts on more specific semantic patterns, training them to enforce the prompted images to be as similar as possible to corresponding natural images. This can preserve more natural semantic patterns as shown in Figure 47 and 8.
Appendix H Stability in Natural Accuracy
As a whole, our method performs more stably in natural accuracy. As shown in Section 4, baselines lose more natural accuracy under many cases, such as the worse transferability and performances under adaptive attacks of C-AVP and the natural accuray drop of Freq shown in Table 3 under adversarially pre-trained models. In comparison, our defense remains high natural accuracy in all of these cases, verifying the stability of our defense.
Appendix I Effectiveness on C&W
C&W method generates adversarial perturbations by performing optimizations in the pixel domain. Differently, our approach additionally considers the frequency domain. It disentangles the frequency domain information and leverages the amplitude and phase spectra as a way to focus more finely on important structural semantics and textures, which are not covered in the compared baselines. Therefore, our method can provide a more effective defense against perturbations generated by C&W.